OpenStudy (anonymous):

If the sine of an acute angle is 3/5 what is the cosine of that angle?

OpenStudy (anonymous):

what is it in fraction form?

OpenStudy (anonymous):

\[(sinx)^2+(cosx)^2=1\]

OpenStudy (anonymous):

so if sin of the angle is 3/5 then just put that into the identity and solve for (cosx)

OpenStudy (anonymous):

it has to be either 4/5 2/5 root2/5 or 1/54

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

ok I will show you :)

OpenStudy (anonymous):

sinx=3/5

OpenStudy (anonymous):

\[(\frac{ 3 }{ 5 })^2+(cosx)^2=1\]

OpenStudy (anonymous):

\[\frac{ 9 }{ 25 }+(cosx)^2=1\]

OpenStudy (anonymous):

\[(cosx)^2=1-(\frac{ 9 }{ 25 })\]

OpenStudy (anonymous):

\[(cosx)^2=\frac{ 16 }{ 25}\]

OpenStudy (anonymous):

then take the square root of both sides!

OpenStudy (anonymous):

(cosx)= ?